### v2

@@ 5 Conclusion
In this paper, we successfully trained a 5-digit 2-layer 3-head addition model that we have verified as known-good. Our investigation revealed that a 6-digit 2-layer 3-head addition model follows the same algorithmic principles, albeit with some implementation variations.
A significant aspect of our work was demonstrating component reuse. We achieved this by integrating an existing addition model into a new “mixed” model designed for both addition and subtraction tasks. This integration allowed us to leverage our in-depth understanding of the addition model’s algorithm to unravel the workings of the mixed model. Our findings indicate that the mixed model employs the nodes of the inserted addition model for conducting both addition and subtraction computations.
This paper lends support to the assertion by Michaud et al. (2023) that many natural prediction problems can be broken down into a finite set of “quanta” computations. These quanta are essential for models to learn in order to minimize loss. Furthermore, it aligns with the concept that comprehending a network’s functionality is fundamentally about identifying and understanding its sub-quanta.
@@ 6 Limitations and Future Work
Our research opens up several promising avenues for future exploration. A key objective is to establish the mixed model as known-good by reverse-engineering its algorithm implementation and validating it through experimental methods. Additionally, we aim to develop a comprehensive model that proficiently handles addition, subtraction, multiplication, and division, ensuring its status as known-good.
Currently, our approach requires significant human effort, which may not be feasible for large-scale applications. However, we believe applying this method to critical tasks, such as logical reasoning involving both simple (e.g., AND, OR) and more complex operators (e.g., Converse, Inverse), is justified. This could lead to the creation of more known-good transformer models, potentially extending to frameworks for ethical decision-making.
A promising finding of our research was the successful insertion of a known-good model into a larger, untrained model while retaining its functionality. We think it might be feasible to embed a compact known-good model into an existing Large Language Model (LLM) to enhance its trustworthiness. Previous studies (Stolfo et al., 2023; Kruthoff, 2024) found nodes in LLMs that were responsible for inaccuracies mathematical calculations. Given our findings, it might be possible to replace these nodes with known-good ones. While this poses an engineering challenge, research by (Voita et al., 2023) and (Hu et al., 2021) indicates the presence of “spare” neurons in LLMs and the potential for integrating small-scale modifications to fix erroneous circuits in LLMs.
@@
A significant challenge in our research was the sensitivity of model behavior to minor changes in the training environment. These changes often altered the specific nodes used for calculations, necessitating manual adjustments to our algorithm-validation test suite. A more effective strategy would be a declarative approach to describing each component in an algorithm hypothesis, allowing automatic identification of corresponding model nodes. For instance, a component description could specify nodes that primarily attend to question digits D2 and D’2, influence answer digit A3, and are crucial for S0 but not S1 complexity questions. Such a declarative language and its accompanying tooling would accelerate research progress and ensure that known-good components retain their functionality when integrated into LLMs. An initial version of this tooling has been created, but significant further work is required to fully realize its potential.
Impact Statement
@@
Acknowledgements
The authors would like to thank Apart Labs (https://apartresearch.com/lab) for their support in producing this paper and more generally supporting the entry of people new to AI Safety into the field.
Appendix A Appendix: Model Configuration
Addition experiments were done in Colab notebooks named “Accurate Addition - Train” and “Accurate Addition - Analyse”. They run on a T4 GPU. Training takes $\sim$15 mins. Analysis takes $\sim$8 mins. The key parameters (which are all configurable) are:
  • n_layers = 2: This is a 2-layer Transformer model
  • n_heads = 3: There are 3 attention heads
  • n_digits = 5: Number of digits in the addition question
Training uses a new batch of data each step (aka Infinite Training Data) to minimise memorisation. During a training run the model processes about 1.5 million training datums. For the 5-digit addition problem there are 100,000 squared (that is 10 billion) possible questions. So the training data is much less than 1% of the possible problems. Because US cascades (e.g. 44445+55555=100000, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare, the data generator was enhanced to increase the likelihood of these cases turning up in the training data
@@ Appendix A Appendix: Model Configuration
Mixed operation (addition and subtraction) experiments were done in Colab notebooks named “Accurate Math - Train” and “Accurate Math - Analyse”. They run on a T4 GPU. Training takes $\sim$25 mins. Analysis takes $\sim$8 mins. The key parameters (which are all configurable) are:
  • n_layers = 3: This is a 3-layer transformer model
  • n_heads = 4: There are 4 attention heads
  • n_digits = 6: Number of digits in the addition and subtraction questions
Training uses a new batch of data each step (aka Infinite Training Data) to minimise memorisation.
@@
In all Colabs, each digit is represented as a separate token. (Liu & Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers”.
Appendix B Appendix: Addition Model Loss
@@
The 1-layer model is not very accurate (see Tab. 5). With 2 layers the model gains accuracy (see Tab. 6). The model uses batch_size = 64, n_heads = 3, lr = 0.00008, weight_decay = 0.1. The loss function is simple:
@@
Table 5: Results for 1-layer 3-head model with 5, 10 & 15 digits
No. digits 5 10 15
Training batches 20K 20K 20K
Loss 0.008314 0.040984 0.031746
Heads used 15 29 46
MLPs used 6 11 16
Table 6: Results for 2-layer 3-head model with 5 & 10 digits
No. digits 5 5 10 10
Training batches 20K 30K 20K 30K
Loss 2.7e-6 2.3e-8 9.4e-5 1.2e-4
Heads used 21 19 57 51
MLPs used 15 15 27 28
The 2-layer, 5-digit model trained over 30 thousand batches has a final training loss of $\sim 2.3\times 10^{-8}$.
Appendix C Appendix: Addition Model Shape Choice
@@
  • Increasing the frequency of hard (cascading US ) examples in the training data so the model has more hard examples to learn from. This improved training speed but did not reduce loss.
@@
The smallest model shape that did reduce loss significantly was 2 layers with 3 attention heads
Appendix D Appendix: TriAdd Implementation
TriAdd transfers data from $(D_{n-1})$ to $(D_{n})$ by integrating the values of $(D_{n}.C)$ and $(D_{n-1}.C)$. This function can be represented as nine bigram mappings with three possible outputs.
Table 7: $D_{n}$.TriAdd can be calculated from $D_{n}.C$ and $D_{n-1}.C$ through nine bigram mappings and yielding the three distinct outputs T8, T9 and T10
$D_{n}.C$ =T8 $D_{n}.C$ =T9 $D_{n}.C$ =T10
$D_{n-1}.C$ = T8 T8 T9 T10
$D_{n-1}.C$ = T9 T8 T9 T10
 $D_{n-1}.C$ = T10 T9 * T10 T10
Note that in the case $D_{n}.C$ = T9 and $D_{n-1}.C$ = T10, the answer is indeterminate. The result could be T8 or T9 but importantly it can not be T10. We choose to use T9 in our definition, but T8 would work just as well.
Appendix E Appendix: Addition Hypothesis 1
@@
  • Is based on the same BA, MC and US operations as the 1-layer.
  • Uses the new early positions to (somehow) do the US calculations with higher accuracy than the 1-layer model.
@@
  • The two attention nodes in the long double staircase positions do the BA and MC calculations and pull in US information calculated in the early positions.
@@
Appendix F Appendix: Addition Hypothesis 2
@@
We claimed the model stores the sum of each digit-pair as a single token in the range “0” to “18” (covering 0+0 to 9+9). We name this operator $D_{n}$.T, where T stands for “token addition”:
  • $D_{n}$.T = $D_{n}$+ $D^{\prime}_{n}$
The $D_{n}$.T operation does not understand mathematical addition. Tab. 8 shows how the model implements the T operator as a bigram mapping.
@@
$D_{n}$.T is a compact way to store data. Tab. 9 show how, if it needs to, the model can convert a $D_{n}$.T value into a one-digit-accuracy BA, MC or MS value.)
Table 9: Converting a $D_{n}$.T value into a BA , MC or MS value.
$D_{n}$.T $D_{n}$.BA $D_{n}$.MC $D_{n}$.MS
@@
Our notation shorthand for one-digit-accuracy these “conversion” bigram mappings is:
  • $D_{n}$.BA = ($D_{n}$.T % 10) where % is the modulus operator
  • $D_{n}$.MC = ($D_{n}$.T // 10) where // is the integer division operator
  • $D_{n}$.MS = ($D_{n}$.T == 9) where == is the equality operator
The D0.T value is perfectly accurate. But the other $D_{n}$.T values are not perfectly accurate because each is constrained to information from just one digit. We define another more accurate operator $D_{n}$.T2 that has “two-digit accuracy”. $D_{n}$.T2 is the pair sum for the nth digit plus the carry bit (if any) from the n-1th digit T:
  • $D_{n}$.T2 = $D_{n}$.T + $D_{n-1}$.MC
$D_{n}$.T2 is more accurate than $D_{n}$.T. The $D_{n}$.T2 value is always in the range “0” to “19” (covering 0+0+0 to 9+9+MakeCarry). Tab. 10 show how the model can implement the T2 operator as a mapping.
Table 10: Calculating $D_{n}$.T2 from $D_{n}$.T and $D_{n-1}$.T
$D_{n}$.T $D_{n}$.T2 if $D_{n-1}$.MC==0 $D_{n}$.T2 if $D_{n-1}$.MC==1
@@
Following this pattern, we define operators $D_{n}$.T3, $D_{n}$.T4 and $D_{n}$.T5 with 3, 4 and 5 digit accuracy respectively:
  • $D_{n}$.T3 = $D_{n}$.T + ( $D_{n-1}$.T2 // 10 ) aka $D_{n}$.T + $D_{n-1}$.MC2
  • $D_{n}$.T4 = $D_{n}$.T + ( $D_{n-1}$.T3 // 10 ) aka $D_{n}$.T + $D_{n-1}$.MC3
  • $D_{n}$.T5 = $D_{n}$.T + ( $D_{n-1}$.T4 // 10 ) aka $D_{n}$.T + $D_{n-1}$.MC4
The value D4.T5 is perfectly accurate as it integrates MC and cascading MS data all the way back to and including D0.T. The values D1.T2, D2.T3, D3.T4 are also all perfectly accurate. If the model knows these values it can calculate answer digits with perfect accuracy:
  • A1 = D1.T2 % 10 with zero loss
  • A2 = D2.T3 % 10 with zero loss
  • A3 = D3.T4 % 10 with zero loss
  • A4 = D4.T5 % 10 with zero loss
  • A5 = D4.T5 // 10 with zero loss
Tab. 11 shows how the hypothesis 2 mathematical framework is applied and Figure 11 shows how this maps to the model. There are nodes that the model needs that this framework does not use, so our hypothesis is not 100% right.
Table 11: Accurately predicting A5 in step 11 and A4 in step 12 using addition hypothesis 2
Calculation Node
D1.T = D1 + D’1
P9.L0.H1 and P9.L0.MLP
D2.T = D2 + D’2
P8.L0.H1 and P8.L0.MLP
D1.T2 = D2.T + (D1.T + (D0 + D0’) // 10) // 10
P10.L0.H1 and P10.L0.MLP
D3.T4 = D3 + D’3 + D2.T3 // 10
P11.L0.H1
D4.T = D4 + D’4
P11.L0.H2
D3.MC3 = D3.T4 // 10
P11.L0.MLP
A5 = (D4.T + D3.MC3) // 10
P11.L1.MLP
D4.T5 = D4.T + D3.T4 // 10
P12.L0.H0
A4 = D4.T5 % 10 // 10
P12.L0.MLP
Figure 11: The 5-digit 2-layer integer addition model algorithm under hypothesis 2
In this hypothesis, all the answer digits are perfectly calculated using the nodes in positions 8 to 11. This hypothesis 2 is feasible, elegant and compact - reflecting the authors (human) values for good code.
Experimenation shows the model does not implement this hypothesis. It retains the long staircase BA calculations in positions 11 to 16. Why? Two reasons suggest themselves:
@@
Appendix G Appendix: Five-Digit Addition Experiments
Figure 12: Our 5-digit integer addition question has 12 question tokens ($D_{n}$,$D^{\prime}_{n}$) followed by 6 answer tokens ($A_{n}$). This known-good 2-layer 3-head model can perform one million predictions without error. The model’s algorithm co-ordinates computation across 9 token positions ($P_{n}$) using 19 attention heads ($H_{n}$) and 15 MLP layers. Three attention heads produce intermediate results that are reused in multiple prediction steps e.g. P10.L0.H1 is reused in 4 prediction steps to calculate answer digits A5 to A2.
The hypothesis 3 pseudo-code was derived iteratively by obtaining experimental results and mapping them to mathematical operations. Figure 12 shows the pseudo-code diagramtically. Some of the experiments and mappings were:
@@ Appendix G Appendix: Five-Digit Addition Experiments
  • Ablating the nodes one by one shows which answer digit(s) are reliant on each node (Ref Figure 7). Most interestingly, ablating P10.L0.H1 impacts the answer digits A5, A4, A3, A2 (but not A1 and A0). This node is used in the calculation of A5, A4, A3, A2 in prediction steps 11, 12, 13 and 14. These relationships are constraints that are all obeyed by the pseudo-code.
  • The pseudo-code has 4 instances where $D_{n}.C$ is calculated using TriCase. PCA of the corresponding nodes (P8.L0.H1, P9.L0.H1, P11.L0.H2 and P14.L0.H1) shows tri-state output for the specified $D_{n}$. (see Figure 13).
@@
  • Activation patching (aka interchange intervention) experiments at attention head level confirmed some aspects of the calculations (see § H for details.
  • The pseudo code includes calculations like D1.C which it says is calculated in P9.L0.H1 and P9.L0.MLP. Ablation tells us both nodes are necessary. For the attention head we use the PCA results for insights. We didn’t implement a similar investigative tool for the MLP layer, so in the pseudo-code we attribute the calculation of D1.C to both nodes.
@@
  • For P15.L0.H1 the attention head PCA could represent either a bi-state or tri-state output (see Figure 14). The pseudo-code shows a bi-state calculation D0.MC at P15.L0.H1, but with small alterations the pseudo-code would work with a tri-state output.
  • The calculation of D1.C2 in P14.L0.H1 is a interesting case. The model needs D1.C2 for A2 accuracy. The model could simply reuse the accurate D1.C2 value calculated in P10. Activation patching shows that it does not. Instead the P14 attention heads calculate D1.C1 from D1 and D’1 directly, and only relies on the P10.D1.C2 value in the case where D1.C2 != D1.C. That is, the calculation is “use P14.D1.C1 value else use P10.D1.C2 values”. This aligns with the model learning the P10.D1.C calculation early in training (for 90% accuracy) and later learning that P10.D1.C2 contains additional information it can use to get to 100% accuracy.

### v7

@@ 5 Conclusion
In this paper, we successfully trained 5-, 6- and 10-digit, 2-layer, 3-head addition models that we have verified as known-good. They all implement the same algorithm, albeit with some implementation variations.
A significant aspect of our work was demonstrating component reuse. We achieved this by integrating an existing addition model into a new larger “mixed” model designed for both addition and subtraction tasks. This mixed model has six nines accuracy.
This integration allowed us to leverage our in-depth understanding of the addition model’s algorithm to unravel the workings of the mixed model’s algorithm. Interestingly, the mixed model re-uses most of the inserted addition model nodes, and “upgrades” most of them to become polysemantic - performing the inserted addition subtask and the corresponding subtraction subtasks simultaneously.
This paper lends support to the assertion by (Michaud et al., 2023) that many natural prediction problems can be broken down into a finite set of “quanta” computations. These quanta are essential for models to learn to minimize loss. Furthermore, it aligns with the concept that comprehending a network’s functionality is fundamentally about identifying and understanding its sub-quanta.
@@ 6 Limitations and Future Work
Our research opens up several promising avenues for future exploration. A key objective is to complete the reverse-engineering of the mixed model’s algorithm. We need to understand how the polysemantic nodes and the OPR+SGN node outputs are represented in the residual stream and show the MLP layers combine this information to select the appropriate question class data to accurately predict the answer digit.
Additionally, we aim to develop a comprehensive model that proficiently handles integer n-digit addition, subtraction, and multiplication, ensuring its status as known-good.
We believe this approach can be re-used to create known-good models in fields like logical reasoning (for AND, OR, Converse, Inverse etc tasks) and planning.
Using a smaller model to improve an existing Large Language Model (LLM) trustworthiness and capabilities is an active area of research. Approaches include composing multiple models (Bansal et al., 2024), LM up-scaling which essentially emulates fine-tuning a large pre-trained model using a small model (Mitchell et al., 2023), and this paper’s insertion of an accurate model into an untrained model. Other research (Voita et al., 2023; Hu et al., 2021) indicates the presence of “spare” neurons in LLMs and the potential for integrating small-scale modifications to fix erroneous circuits in LLMs. Further research into the benefits of using known-good models to increase LLM trustworthiness and capabilities is justified.
@@
Our automated framework, for discovering the location of algorithm sub-tasks in models, sped up our research. Our framework can be re-used to ensure that known-good components retain their functionality when integrated into LLMs.
Some of our automated framework is specific to our maths models - much is generic and reuseable on other types of models. Our automation focuses more on attention heads and less on MLP layers. Extending our framework to integrate the output data from complimentary frameworks (e.g. ACDC’s node connectivity strength data (Conmy et al., 2023)) would provide more automated insights into a model’s behavior, further speeding up research.
We made progress on a declarative language to describe an algorithm in terms of the number and types sub-tasks that must exist in a model for the algorithm to feasibly explain the model’s implementation. We made progress on a framework to automatically test an algorithm described in this way against a specific model. Further work is required to fully realize its potential.
7 Impact Statement
@@
Appendix A Appendix: Terminology
These terms and abbreviations are used in this paper and the associated Colabs and python code:
  • Pn : Model (input or output) token position. Zero-based. e.g. P18, P18L1H0
  • Ln : Model layer n. Zero-based. e.g. P18L1H2
  • Hn : Attention head n. Zero-based. e.g. P18L1H2
  • Mn : MLP neuron n. Zero-based
  • PnLnHn : Location / name of a single attention head, at a specified layer, at a specific token position
  • PnLnMn : Location / name of a single MLP neuron, at a specified layer, at a specific token position
  • D : First number of the pair question numbers
  • Dn : nth numeric token in the first question number. Zero-based. D0 is the units value
  • D’ : Second number of the pair question numbers
  • D’n : nth token in the second question number. Zero-based. D0 is the units value
  • A : Answer to the question (including answer sign)
  • An : nth token in the answer. Zero-based. A0 is the units value. The highest token is the “+" or "-" answer sign
  • S : Prefix for Addition. Think S for Sum. Aka ADD.
  • SA : Basic Add. An addition sub-task. An.SA is defined as (Dn + D’n) % 10. e.g. 5 + 7 gives 2
  • SC : Carry One. An addition sub-task. An.SC is defined as Dn + D’n >=>=> = 10. e.g. 5 + 7 gives True
  • SS : Make Sum 9. An addition sub-task. An.SS is defined as Dn + D’n == 9. e.g. 5 + 7 gives False
  • ST : TriCase. An addition sub-task. Refer paper 2 for details
  • ST8, ST9, ST10 : Outputs of the ST TriCase sub-task.
  • M : Prefix for Subtraction with a positive answer. Think M for Minus. Aka SUB
  • : Basic Difference. A subtraction sub-task. An.MD is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • : Borrow One. A positive-answer subtraction sub-task. An.MB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
  • MZ : Make Zero. A positive-answer subtraction sub-task. An.MZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
  • MT : TriCase. A positive-answer subtraction sub-task.
  • MT1, MT0, MT-1 : Outputs of the MT TriCase sub-task.
  • N : Prefix for Subtraction with a negative answer. Think N for Negative. Aka NEG
  • ND : Basic Difference. A negative-answer subtraction sub-task. An.ND is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • NB : Borrow One. A negative-answer subtraction sub-task. An.NB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
  • NZ : Make Zero. A negative-answer subtraction sub-task. An.NZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
  • NT : TriCase. A negative-answer subtraction sub-task.
  • GT : Greater Than. A (positive-answer or negative-answer) subtraction sub-task. An.GT is defined as Dn $>$ D’n. e.g. 3 $>$ 5 gives False
  • OPR : Operator. A sub-task that attends to the + or - token in the question (which determines whether the question is addition or subtraction).
  • SGN : Sign. A sub-task that attends to the first answer token, which is + or -
  • PCA : Principal Component Analysis
  • EVR : Explained Variance Ratio. In PCA, EVR represents the percentage of variance explained by each of the selected components.
Appendix B Appendix: Model Configuration
@@
Addition, subtraction and mixed (addition and subtraction) training experiments were done in a Colab notebook. The Colab runs on a T4 GPU. Training takes up to 35 mins. The key parameters (and their common configurations) are:
  • n_layers = 1, 2 or 3: Number of layers.
  • n_heads = 3 or 4: Number of attention heads.
  • n_digits = 5, 6 or 10: Number of digits in the question.
@@
Each digit is represented as a separate token. (Liu and Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers". The model’s vocabulary contains 14 tokens ( 0, .., 9, +, -, =, *, / ) to enable this and planned future investigations.
Training uses a new batch of data each step (aka Infinite Training Data) to minimise memorisation. Depending on the configuration, each training run processes 1 to 4 million training datums. For the 5-digit addition problem there are 100,000 squared (that is 10 billion) possible questions. So the training data is much less than 1% of the possible problems.
Addition and subtraction include rare edge cases. For example, the SS cascades (e.g. 44445+55555=100000, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare. The data generator was enhanced to increase the frequency of all known edges cases. This lead to lower model loss.
Validation test data covering all edge cases was manually constructed. These test cases are not used during training.
Appendix C Appendix: Model Loss
@@
The model defaults to batch size = 64, learning rate = 0.00008 and weight decay = 0.1. The loss function is simple:
@@
In our experimental models, the number of digits in the question varies from 5 to 10, the number of layers varies from 1 to 4, the number of heads varies from 3 to 4. Each experimental model’s loss is detailed in Tab. 6.
Appendix D Appendix: Addition Model Shape
@@
  • Increasing the frequency of hard (cascading SS ) examples in the training data so the model has more hard examples to learn from. This improved training speed but did not reduce loss.
@@
The smallest model shape that did reduce loss significantly was 2 layers with 3 attention heads.
Appendix E Appendix: Experimental models
Twenty-one experiment models were trained and analyzed (refer Tab. 6). The models and output of analysis are available on HuggingFace at https://huggingface.co/PhilipQuirke/VerifiedArithmetic
Num Digits Num Layers Num Heads Training Steps Seed Addition / Sub Accuracy Training loss Heads used MLPs used
Addition models
5 1 3 30K 372001 two nines 9.4e-2 15 6
5 2 3 15K 372001 six nines 1.6e-8 30 16
5 2 3 40K 372001 six nines 2.0e-9 22 15
6 2 3 15K 372001 six nines 1.7e-8 31 17
6 2 3 20K 173289 six nines 1.5e-8 28 17
6 2 3 20K 572091 six nines 7.0e-9 35 17
6 2 3 40K 372001 six nines 2.0e-9 29 17
10 2 3 20K 572091 six nines 7.0e-9 44 28
Subtraction models
6 2 3 30K 372001 six nines 5.8e-6 40 21
10 2 3 75K 173289 two nines 2.0e-3 101 37
Mixed models
6 3 4 40K 372001 six / six nines 5.0e-9 54 26
10 3 4 75K 173289 five / two nines 1.1e-6 143 53
Mixed models initialized with addition model
6 2 3 40K 572091 six / five nines 2.4e-8 57 21
6 3 3 40K 572091 six / five nines 1.8e-8 70 35
6 3 3 80K 572091 six / five nines 1.6e-8 75 35
6 3 4 40K 372001 six / six nines 8.0e-9 72 26
6 3 4 40K 173289 five / five nines 1.4e-8 60 29
6 3 4 50K 572091 six / five nines 2.9e-8 79 29
10 3 3 50K 572091 five / five nines 6.3e-7 90 45
Mixed models initialized with add model. Reset useful heads every 100 steps
6 4 4 40K 372001 five / five nines 1.7e-8 51 30
Mixed models initialized with add model. Reset useful heads & MLPs every 100 steps
6 4 3 40K 372001 four / two nines 3.0e-4 115 53
Table 6: Main experimental models studied
For each model the ’VerifiedArithmeticTrain’ Colab notebook generates two files:
  • A “XXXXXX.pth" file containing the model weights
  • A “XXXXXX_train.json" file containing configuration information and training loss data
While, for each model the ’VerifiedArithmeticAnalysis’ Colab notebook generates two more files:
  • A “XXXXXX_behavior.json" file containing generic “behavior" facts learnt about the model by the Colab e.g. P18L0H0 attends to tokens D3 and D’3
  • A “XXXXXX_maths.json" file containing “maths-specific" facts learnt about the model by the Colab e.g. P18L0H0 performs the A3.SC sub-task.
Appendix F Appendix: TriAdd Implementation
TriAdd transfers data from $A_{n-1}$to $A_{n}$by integrating the values of $A_{n-1}.ST$ and $A_{n}.ST$ This function can be represented as nine bigram mappings with three possible outputs. (Refer Tab.7.)
$A_{n}.ST$ $A_{n}.ST$ $A_{n}.ST$
$A_{n-1}.ST$ = ST8 = ST9 = ST10
ST8 ST8 ST9 ST10
ST9 ST8 ST9 ST10
 ST10 ST8 * ST10 ST10
Table 7: $A_{n}$.TriAdd can be calculated from $A_{n}.ST$ and $A_{n-1}.ST$ through nine bigram mappings and yielding the three distinct outputs ST8, ST9 and ST10
Note that in the case $A_{n}.ST$ = ST9 and $A_{n-1}.ST$ = ST10, the answer is indeterminate. The result could be ST8 or ST9 but importantly it can not be ST10. We choose to use ST8 in our definition, but ST9 would work just as well.
Appendix G Appendix: Complexity
To analyze question difficulty, we categorized addition questions by the complexity of the computation required to solve the question, as shown in Tab. 8. The categories are arranged according to the number of digits that a carry bit has to cascade through.
Name Contains Example Freq
S0 SA 11111+12345=23456 $\sim$5%
S1 SA,SC 11111+9=22230 $\sim$21%
S2 SA,SCx2 11111+89=22300 $\sim$34%
S3 SA,SCx3 11111+889=23000 $\sim$28%
S4 SA,SCx4 11111+8889=30000 $\sim$11%
S5 SA,SCx5 11111+88889=100000 $\sim$2%
Table 8: We categorise addition questions into non-overlapping “calculation complexity" quanta, ordered by increased computational difficulty (and decreasing occurrence frequency). Five-digit addition questions quanta are S0 to S5. Ten-digit addition question quanta are S0 to S10. S10’s frequency is $\sim 3e-4$ showing the need to enrich training data for rare edge cases.
Appendix H Appendix: Addition Hypothesis 1
@@
  • Is based on the same SA, SC and SS operations as the 1-layer.
  • Uses the new early positions to (somehow) do the SS calculations with higher accuracy than the 1-layer model.
@@
  • The two attention nodes in the long double staircase positions do the SA and SC calculations and pull in SS information calculated in the early positions.
@@
Appendix I Appendix: Addition Hypothesis 2
@@
We claimed the model stores the sum of each digit-pair as a single token in the range “0” to “18” (covering 0+0 to 9+9). We name this operator $A_{n}$.T, where T stands for “token addition”:
  • $A_{n}$.T = $D_{n}$+ $D^{\prime}_{n}$
The $A_{n}$.T operation does not understand mathematical addition. Tab. 9 shows how the model implements the T operator as a bigram mapping.
@@
$A_{n}$.T is a compact way to store data. Tab. 10 show how, if it needs to, the model can convert a $A_{n}$.T value into a one-digit-accuracy SA, SC or SS value.)
$A_{n}$.T $A_{n}$.SA $A_{n}$.SC $A_{n}$.SS
@@
Table 10: Converting a $A_{n}$.T value into a SA , SC or SS value.
Our notation shorthand for one-digit-accuracy these “conversion" bigram mappings is:
  • $A_{n}$.SA = ($A_{n}$.T % 10) where % is the modulus operator
  • $A_{n}$.SC = ($A_{n}$.T // 10) where // is the integer division operator
  • $A_{n}$.SS= ($A_{n}$.T == 9) where == is the equality operator
The A0.T value is accurate. But the other $A_{n}$.T values are not accurate because each is constrained to information from just one digit. We define another more accurate operator $A_{n}$.T2 that has “two-digit accuracy”. $A_{n}$.T2 is the pair sum for the nth digit plus the carry bit (if any) from the n-1th digit T:
  • $A_{n}$.T2 = $A_{n}$.T + $A_{n-1}$.SC
$A_{n}$.T2 is more accurate than $A_{n}$.T. The $A_{n}$.T2 value is always in the range “0” to “19” (covering 0+0+0 to 9+9+CarryOne). Tab. 11 show how the model can implement the T2 operator as a mapping.
$A_{n}$.T $A_{n}$.T2 if $A_{n-1}$.SC==0 $A_{n}$.T2 if $A_{n-1}$.SC==1
@@
Table 11: Calculating $A_{n}$.T2 from $A_{n}$.T and $A_{n-1}$.T
Following this pattern, we define operators $A_{n}$.T3, $A_{n}$.T4 and $A_{n}$.T5 with 3, 4 and 5 digit accuracy respectively:
  • $A_{n}$.T3 = $A_{n}$.T + ( $A_{n-1}$.T2 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC2
  • $A_{n}$.T4 = $A_{n}$.T + ( $A_{n-1}$.T3 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC3
  • $A_{n}$.T5 = $A_{n}$.T + ( $A_{n-1}$.T4 // 10 ) aka $A_{n}$.T + $A_{n-1}$.SC4
The value A4.T5 is accurate as it integrates SC and cascading SS data all the way back to and including A0.T. The values A1.T2, A2.T3, A3.T4 are also all accurate. If the model knows these values it can calculate answer digits accurately:
  • A1 = A1.T2 % 10
  • A2 = A2.T3 % 10
  • A3 = A3.T4 % 10
  • A4 = A4.T5 % 10
  • A5 = A4.T5 // 10
In this hypothesis, all the answer digits are accurately calculated using the nodes in positions 8 to 11. This hypothesis 2 is feasible, elegant and compact - reflecting the authors (human) values for good code.
Experimenation shows the model does not implement this hypothesis. It retains the long staircase SA calculations in positions 11 to 16. Why? Two reasons suggest themselves:
@@
Appendix J Appendix: Addition Hypothesis 3
The hypothesis 3 pseudo-code was derived iteratively by obtaining experimental results and mapping them to mathematical operations. Some of the experiments and mappings were:
@@
  • Ablating the nodes one by one shows which answer digit(s) are reliant on each node (Ref Table 1). Most interestingly, ablating P10.L0.H1 impacts the answer digits A5, A4, A3, A2 (but not A1 and A0). This node is used in the calculation of A5, A4, A3, A2 in prediction steps 11, 12, 13 and 14. These relationships are constraints that are all obeyed by the pseudo-code.
  • The pseudo-code has 4 instances where $A_{n}.ST$ is calculated using TriCase. PCA of the corresponding nodes (P8.L0.H1, P9.L0.H1, P11.L0.H2 and P14.L0.H1) shows tri-state output for the specified $D_{n}$. (see Figure 8).
@@
  • Activation patching (aka interchange intervention) experiments at attention head level confirmed some aspects of the calculations (see § K for details.
  • The pseudo code includes calculations like A1.ST which it says is calculated in P9.L0.H1 and P9.L0.MLP. Ablation tells us both nodes are necessary. For the attention head we use the PCA results for insights. We didn’t implement a similar investigative tool for the MLP layer, so in the pseudo-code we attribute the calculation of A1.ST to both nodes.
@@
  • For P15.L0.H1 the attention head PCA could represent either a bi-state or tri-state output (see Figure 9). The pseudo-code shows a bi-state calculation A0.SC at P15.L0.H1, but with small alterations the pseudo-code would work with a tri-state output.
  • The calculation of A1.ST2 in P14.L0.H1 is a interesting case. The model needs A1.ST2 for A2 accuracy. The model could simply reuse the accurate A1.ST2 value calculated in P10. Activation patching shows that it does not. Instead the P14 attention heads calculate A1.ST1 from D1 and D’1 directly, and only relies on the P10.D1.ST2 value in the case where A1.ST2 != A1.ST. That is, the calculation is “use P14.A1.ST1 value else use A1.ST2 values". This aligns with the model learning the P10.A1.ST calculation early in training (for 90% accuracy) and later learning that P10.A1.ST2 contains additional information it can use to get to six nines accuracy.
Figure 5: For a sample 5-digit 2-layer 3-head addition model, this map shows a compacted view of all useful token positions (horizontally) and all useful attention heads and MLP layers (vertically) used in predictions as green cells. Each cell shows the simpliest (lowest complexity) quanta S0, S1, etc impacted when we ablate each node. To answer S0 questions, only the S0 nodes are used. To answer S1 questions, S0 and S1 nodes are used, etc. The model only uses nodes in nine token positions.
Figure 6: This map shows the % of enriched questions that fail when we ablate each node in a 5-digit 2-layer 3-head addition model. The model only uses nodes in token positions P8 to P16 (i.e. tokens D’2 to A1). Lower percentages correspond to rarer edge cases. The grey space represents nodes that are not used by the model.
Figure 7: This map shows the input tokens each attention head attends to at each token position in a 5-digit 2-layer 3-head addition model. At token position P11 the model predicts the first answer digit A5. All digit pairs (e.g. D2 D’2) are attended to by P11.